Identify solutions to systems of equations:
-2x-4y=1
12y=-6x-3

Answer by ewatrrr(24785)

3(-2x-4y=1)
6x + 12y=-3
0 = 0
Infinite solutions
12y=-6x-3 0r 4y = -2x + 1
All Ordered Pairs: (x, (1-2x)/4)are the Solutions